Shift report 20000724 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Mon Jul 24 2000 - 00:04:48 EDT

  • Next message: http account: "Shift report 20000724 00:00-08:00"

    ------------------------------------------------------------
    Date: Mon Jul 24 00:04:48 2000, Shift: 16:00-24:00
    Supervisor: ykl, with: yb
    Log book#3 pages 60 - 60
    
    Shift Summary:
    ------------------------------------------------------------
    1.Some PS problem and RF pproblems. There was two and
    half hours of stable beams at 65G although yellow beam was
    rather low. The beams were steered to PHENIX.
    2. Runs at 65G, cogged; Run number 1900 and 1904. Collision
    rates were low.
    3. The HV controls for C1, H1, etc, still stays
    incommunicado. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Mon Jul 24 2000 - 00:15:41 EDT